# Oligomorphic groups, categories of partial bijections, and ultrahomogeneous cubic spaces over finite fields

### v2

@@
We show that for certain class of oligomorphic groups there is a version of multiplication of double cosets in the Ismagilov–Olshanski sense. Categories of (reduced) double cosets are realized as certain categories of partial bijections. As an example, we consider the ultrahomogeneous linear space $\mathcal{Z}$ of countable dimension over a finite field equipped with a cubic form. The group of automorphisms of $\mathcal{Z}$ is an oligomorphic group, we describe its open subgroups. According Tsankov, this gives a classification of its unitary representations. The category of reduced double cosets in this case is the category of partial linear bijections of finite-dimensional cubic spaces preserving cubic forms.
@@
0.1. Categories of double cosets. In theory of unitary representations of ’large’ (’infinite-dimensional’ groups [^2] A. M. Vershik proposed the term ’large groups’ instead of ’infinite-dimensional groups’. Numerous ’infinite-dimensional groups’ are totally disconnected, so they have dimension 0 as topological spaces. This is so for groups of types $2^{\circ}$ and $4^{\circ}$-$7^{\circ}$ below. We emphasize that infinite dimensional unitary groups and (zero-dimensional) infinite symmetric groups are close relatives (at least from the point of view of representation theory).), quite often the following phenomena arise. Let $G$ be a group, and $K$ is a subgroup. It can happen that the space of double cosets $K\backslash G/K$ admits a natural associative multiplication
@@
Remarks. a) In all cases, $K$ looks as an infinite-dimensional counterpart of compact groups. In [25], Chapter VIII, there was proposed a heuristic notion of a ’heavy group’. It turns out to be that for groups $G$ with a heavy subgroup $K$ the multiplication of double cosets arises more or less automatically [^3] Examples of heavy groups are the complete symmetric group $S(\infty)$, the complete infinite hyperoctahedral group, the complete infinite-dimensional unitary group $\mathrm{U}(\infty)$, the complete orthogonal group $\mathrm{O}(\infty)$, the group of measure preserving transformations of a Lebesgue measure space. Oligomorphic groups discussed below look like ’heavy groups’.
@@
This statement allows to easily describe multiplications of reduced double cosets for oligomorphic groups (under some aditional conditions), the corresponding categories are certain categories of partial bijections (the term ’reduced double cosets’ is explained below in Subsect. 1).
The work [40] was used earlier in [29], [30]. Since oligomorphic groups look like heavy groups, the next (open) question is a search of interesting overgroups of oligomorphic groups, see a discussion in Subsect. 2.
@@ 1. Multiplications of double cosets and trains of ’large’ groups
We denote classes of equivalence (points of the quotient) as $K_{1}gK_{2}$.
@@
Equivalently, let us consider the group algebra ${\mathbb{C}}[G]$ of $G$, denote $\varphi\star\psi$ the convolution in ${\mathbb{C}}[G]$. For $p\in G$ denote by $\delta_{p}\in{\mathbb{C}}[G]$ the function, which is 1 at $p$ and zero at all other points, so $\delta_{p}\star\delta_{q}=\delta_{pq}$.
@@
$\varphi(k_{1}gk_{2})=\varphi(g)\qquad\text{for any $g\in G$, $k_{1}$, $k_{2}\in K$.}$
@@
Remark. Olshanski [35], [37], [38] proposed the following explanation of this phenomenon. Let $G_{n}$ be finite-dimensional (or finite) groups, let $G_{\infty}$ be their ’limit’ (as symmetric groups $S_{n}$ and $S_{\infty}$). Then, in formulas of type (1.2), we get a concentration of convolutions near one double coset (but this explanation does not always works; moreover, generally, this does not formally imply (1.4)). $\boxtimes$
@@
Consider a separable topological group $G$, its subgroup $K$, and a family of subgroups $K_{\alpha}\subset K$ enumerated by a set $\mathcal{A}$ (and this family includes $K$). For any unitary representation $\rho$ of $G$ in a Hilbert apace $\mathcal{L}$, we denote by $\mathcal{L}_{\alpha}$ the subspace of $K_{\alpha}$-fixed vectors, by $P_{\alpha}$ the orthogonal projection to $\mathcal{L}_{\alpha}$. Consider operators
@@
Let $\mathcal{K}$ be a category. Denote by $\operatorname{Ob}(\mathcal{K})$ the set of objects of $\mathcal{K}$, by $\operatorname{Mor}(V,W)=\operatorname{Mor}_{\vphantom{|^{|}}\mathcal{K}}(V,W)$ the set all morphisms $V\to W$, by $\operatorname{End}(V)$ the semigroup of all endomorphisms of an object $V$. We assume that these semigroups have units $1_{V}$ and these units are units in the category $\mathcal{K}$, i.e., for any $\mathfrak{p}\in\operatorname{Mor}(V,W)$ we have
@@
A *-representation $(Q,\varkappa)$ of a category $\mathcal{K}$ is a functor from $\mathcal{K}$ to the category of Hilbert spaces and bounded operators. Namely, for each $V\in\operatorname{Ob}(\mathcal{K})$ we assign a Hilbert space $Q(V)$ and for each $\mathfrak{p}\in\operatorname{Mor}(V,W)$ we assign a bounded operator
@@
Arthur Lieberman [20] obtained the classification of all unitary representations of $S_{\infty}$. Any irreducible representation of this group is induced from an irreducible representation of some subgroup $K^{\circ}_{\alpha}$ trivial on $K_{\alpha}$.
@@
Recall the definition of partial bijections (see, e.g., [25], Sect. VIII.1). Let $X$, $Y$ be sets. A partial bijection $\mu:X\to Y$ is a bijection from a subset $A\subset X$ to a subset $B\subset Y$. We say that $A$ is the domain $\operatorname{dom}\mu$ of $\mu$, $B$ is the image $\operatorname{im}\mu$. A product of partial bijections $\mu:X\to Y$, $\nu:Y\to Z$ is defined in the following way: $\operatorname{dom}\nu\mu:=\mu^{-1}(\operatorname{dom}\nu\,\cap\,\operatorname{im}\mu)$ and $\nu\mu(x)=\nu(\mu(x))$ for $x\in\operatorname{dom}\nu\mu$. We also define the pseudoinverse partial bijection, if $\mu:X\to Y$ is as above, then its pseudoinverse $\mu^{*}:X\to Y$ is the inverse bijection $B\to A$.
Denote by ${\operatorname{PB}}$ the category of partial bijections of finite sets. According [35] (see also exposition in [25], Sect. VIII.1-2), the following theorem holds:
@@
a) The category ${\operatorname{PB}}$ is the train $\operatorname{Tra}\bigl{(}S_{\infty},\{K_{\alpha}\}\bigr{)}$.
b) The functor $\rho\mapsto\widetilde{\rho}_{\alpha,\beta}$ is a one-to-one correspondence between unitary representations of $S_{\infty}$ and *-representations of the category ${\operatorname{PB}}$.
Classification of *-representations of the category ${\operatorname{PB}}$ is a simple problem, see [25], Sect. VIII.1-2 (also, see below, Subsect. 2). In fact, Theorem 1.1 is equivalent to Lieberman’s description [20] of unitary representations of $S_{\infty}$.
@@ 2. Oligomorphic groups and their trains
b) For any irreducible representation $\rho$ of $G$ there is an open subgroup $H^{\circ}$ and a subgroup $H\subset H^{\circ}$ of finite index such that $\rho$ is inducedl [^7] For induced representations of totally disconnected groups, see Mackey [21], [23]. Also, see below Subsect. 2. from a representation of $H^{\circ}$ trivial on $H$.
@@
$\Xi:=\Omega\setminus B,\quad G:=K(B),\quad M:=p(A)\setminus B,\quad N:=q^{-1}B\setminus C$
@@
Remark. So $\operatorname{Tra}(K)$ is realized by partial bijection and the set of morphisms is closed with respect to the involution. Therefore, it is an ’inverse category’ in the sense of [16], i.e., a categorical counterpart of inverse semigroups. $\boxtimes$
@@
b) The map $\rho\mapsto\widetilde{\rho}$ is a bijection between unitary representations of $K$ and *-representations of the category $\operatorname{Tra}(K)$.
@@
The proof of this statement given below in Subsect. 22 is a simple corollary of Tsankov’s Theorem 2.1, formal manipulations with ’ordered categories’ in the sense of [25], and Mackey’s technique of induced representations [21][23], [2].
@@
If $\iota(A)\subset B$, then $\iota$ is fixed by the subgroup $K(B)$, and this does not impose additional conditions for $F$. $\square$
@@
b) Additionally, let $(Q,\varkappa)$ be irreducible. Then the $\preceq$-minimal perfect set such that $Q(D)\neq 0$ is unique (upto $K$-equivalence).
We call this $\preceq$-minimal perfect set $D$ the conductor of an irreducible representation.
@@
Let $(Q,\varkappa)$ be a representation of a category $\mathcal{K}$. Let $C\in\operatorname{Ob}(\mathcal{K})$. Let $\xi_{\alpha}$ be a family of vectors in $Q(C)$. For any $D\in\operatorname{Ob}(\mathcal{K})$ we consider the closed linear span $N(D)$ of all vectors $\varkappa_{\vphantom{|^{|}}D,C}(\mu)\xi_{\alpha}$, where $\mu$ ranges in $\operatorname{Mor}(C,D)$. Clearly, the family of subspaces $N(D)$ is a subrepresentation (we call it the cyclic span of $\xi_{\alpha}$). We say that a family $\xi_{\alpha}$ is cyclic if $N(D)=Q(D)$ for all $D$.
@@
A *-representation $(Q,\varkappa)$ of $\operatorname{Tra}(K)$ with a cyclic space $Q(C)$ is uniquely determined by the representation of $\operatorname{End}(C)$ in $Q(C)$.
Proof. This is a version of the ’GNS-construction’. Let $v_{1}$, $v_{2}\in Q(C)$, $r_{1}$, $r_{2}\in\operatorname{Mor}(C,D)$. Then
@@
Proof. Consider a sequence $A_{1}$, $A_{2}$, $\dots$ including all perfect sets in $\Omega$ defined up to the $K$-equivalence. Since $K$ is oligomorphic, it has only finite number of orbits on the space of finite subsets of a fixed cardinality. So, we can assume that numbers of elements of $A_{j}$ are (non-strictly) increasing. We take the cyclic span $(Q^{(1)},\varkappa^{(1)})$ of $R(A_{1})$ and the orthogonal complement $(R^{(1)},\rho^{(1)})$ of the cyclic span. By construction, we have $R^{(1)}(A_{1})=0$. Next, we take the cyclic span $(Q^{(2)},\varkappa^{(2)})$ of $R^{(1)}(A_{1})$ and orthogonal complement $(R^{(2)},\rho^{(2)})$ of $(Q^{(2)},\varkappa^{(2)})$. Now we have $R^{(2)}(A_{1})=0$, $R^{(2)}(A_{2})=0$. Etc.
@@
Proof. It is sufficient to verify the statement for irreducible representations. Consider an exhausting family of subsets $B\subset R_{1}\subset R_{2}\subset\dots\subset\Omega\setminus B$. By Lemma 2.4, for each $j$ there exists $g_{j}\in K(B)$ such that
@@
Consider the quasiregular representation $\pi$ of $K$ in $\ell^{2}(\mathcal{E}(A,\Omega))$, see (2.1). For $\iota\in\mathcal{E}(A,\Omega)$ denote by $e_{\iota}$ the $\delta$-function supported by $\iota$. If $\iota\in\mathcal{E}(A,B)$, then $\iota$ is fixed by all $g_{j}$. Let $\iota(A)\not\subset B$. Fix $\nu\in\mathcal{E}(A)$. There is $N$ such that $(\iota(A)\setminus B)$, $(\nu(A)\setminus B)\subset R_{N}$. For large $j\geqslant N$ we have
@@
2.7. Examples. We discuss an initial collection of examples of oligomorphic groups (for $1^{\lozenge}$-$4^{\lozenge}$, see [40], Sect. 5, and the corresponding categories.
$0^{\lozenge}.$ Infinite symmetric group, see our Subsect. 1.
$1^{\lozenge}.$ Consider the group $G=\operatorname{Aut}({\mathbb{Q}},<)$ of all order preserving bijections of the set ${\mathbb{Q}}$ of rational numbers. Perfect sets are arbitrary finite subsets $A\subset{\mathbb{Q}}$. We come to the category $\operatorname{Tra}\bigl{(}\operatorname{Aut}({\mathbb{Q}},<)\bigr{)}$. It objects are finite ordered sets. Morphisms $A\to B$ are order preserving partial bijections. Groups of automorphisms are trivial.
So, irreducible unitary representations are enumerated by the number #$\#A$ of elements of a conductor $A$. In other words, any irreducible representations of this group is realized in the spaces $\ell^{2}$ on the set of all $k$-element subsets in ${\mathbb{Q}}$ for some $k$.
$2^{\lozenge}.$ Let $\mathcal{R}$ be the Rado graph, see, e.g., [8]. Recall that it is a graph with a countable number of vertices satisfying the condition (the ’Urysohn extension property’): let $A\subset B$ be finite graphs, then any embedding $A\to\mathcal{R}$ extends to an embedding $B\to\mathcal{R}$.
Consider a countable set, call its elements ’vertices’, for each pair of vertices draw an edge connecting them with probability $1/2$. Then we get the Rado graph with probability 1, see [10], [9].
The group $\operatorname{Aut}(\mathcal{R})$ of automorphisms of $\mathcal{R}$ acting on the set of vertices is perfect oligomorphic, arbitrary finite subset of the set of vertices is perfect. Objects of the train $\operatorname{Tra}(\operatorname{Aut}(\mathcal{R}))$ are finite graphs, a morphism $X\to Y$ is an isomorphism of an induced subgraph [^9] To avoid a misleading we use the term ’induced subgraph’: we consider some set of vertices of $X$ and all edges of $X$ connecting these vertices. $A\subset X$ to an induced subgraph $B\subset Y$.
@@
$3^{\lozenge}.$ Consider a linear space $V$ over a finite field $\mathbb{F}$ consisting of sequences $v=(v_{1},v_{2},\dots)$, where $v_{j}\in\mathbb{F}$ and $v_{k}=0$ for sufficiently large $k$. The group $\operatorname{GL}(\infty,\mathbb{F})$ of all linear transformations of $V$ can be regarded as the group of all invertible infinite matrices over $\mathbb{F}$, which have finite number of elements in each column. The group $\operatorname{GL}(\infty,\mathbb{F})$ is perfect oligomorphic, perfect subsets are finite-dimensional linear subspaces in $V$.
@@
Objects of the train are finite dimensional linear spaces over $\mathbb{F}$. A morphism $X\to Y$ is a partial linear bijection, i.e., an invertible operator from a subspace $A\subset X$ to a subspace $B\subset Y$. Groups of automorphisms are group of linear transformations $\operatorname{GL}(X)$. So irreducible representation are enumerated by a number $n=0$, $1$, $2$, $\dots$ and a complex irreducible representation of the group $\operatorname{GL}(n,\mathbb{F})$.

### v3

@@
We show that for a certain class of oligomorphic groups there is a version of multiplication of double cosets in the Ismagilov–Olshanski sense. Categories of (reduced) double cosets are realized as certain categories of partial bijections. As an example, we consider the ultrahomogeneous linear space $\mathcal{Z}$ of countable dimension over a finite field equipped with a cubic form. The group of automorphisms of $\mathcal{Z}$ is an oligomorphic group, we describe its open subgroups. According Tsankov, this gives a classification of its unitary representations. The category of reduced double cosets in this case is the category of partial linear bijections of finite-dimensional cubic spaces preserving cubic forms.
@@
0.1. Categories of double cosets. In theory of unitary representations of ’large’ (’infinite-dimensional’ groups [^2] A. M. Vershik proposed the term ’large groups’ instead of ’infinite-dimensional groups’. Numerous ’infinite-dimensional groups’ are totally disconnected, hence they have dimension 0 as topological spaces. This is so for groups of types $2^{\circ}$ and $4^{\circ}$-$7^{\circ}$ below. We emphasize that infinite dimensional unitary groups and (zero-dimensional) infinite symmetric groups are close relatives (at least from the point of view of representation theory).), quite often the following phenomena arise. Let $G$ be a group, and $K$ is a subgroup. It can happen that the space of double cosets $K\backslash G/K$ admits a natural associative multiplication
@@
Remarks. a) In all cases, $K$ looks as an infinite-dimensional counterpart of compact groups. In [26], Chapter VIII, there was proposed a heuristic notion of a ’heavy group’. It turns out that for groups $G$ with a heavy subgroup $K$ the multiplication of double cosets arises more or less automatically [^3] Examples of heavy groups are the complete symmetric group $S(\infty)$, the complete infinite hyperoctahedral group, the complete infinite-dimensional unitary group $\mathrm{U}(\infty)$, the complete orthogonal group $\mathrm{O}(\infty)$, the group of measure preserving transformations of a Lebesgue measure space. Oligomorphic groups discussed below look like ’heavy groups’.
@@
This statement allows us to easily describe multiplications of reduced double cosets for oligomorphic groups (under some additional conditions). The corresponding categories are certain categories of partial bijections (the term ’reduced double cosets’ is explained below in Subsect. 1).
The work [41] was used earlier in [30], [31]. Since oligomorphic groups look like heavy groups, the next (open) question is a search for interesting overgroups of oligomorphic groups, see a discussion in Subsect. 2.
@@ 1. Multiplications of double cosets and trains of ’large’ groups
We denote classes of equivalence (points of the quotient) by $K_{1}gK_{2}$.
@@
Equivalently, let us consider the group algebra ${\mathbb{C}}[G]$ of $G$, denote by $\varphi\star\psi$ the convolution in ${\mathbb{C}}[G]$. For $p\in G$ denote by $\delta_{p}\in{\mathbb{C}}[G]$ the function, which is 1 at $p$ and zero at all other points. So, $\delta_{p}\star\delta_{q}=\delta_{pq}$.
@@
$\varphi(k_{1}gk_{2})=\varphi(g)\qquad\text{for any $g\in G$, and $k_{1}$, $k_{2}\in K$.}$
@@
Remark. Olshanski [36], [38], [39] proposed the following explanation of this phenomenon. Let $G_{n}$ be finite-dimensional (or finite) groups, let $G_{\infty}$ be their ’limit’ (as symmetric groups $S_{n}$ and $S_{\infty}$). Then, in formulas of type (1.2), we get a concentration of convolutions near one double coset (but this explanation does not always works; moreover; also, this does not formally implies (1.4)). $\boxtimes$
@@
Consider a separable topological group $G$, its subgroup $K$, and a family of subgroups $K_{\alpha}\subset K$ enumerated by a set $\mathcal{A}$ (we assume that this family includes $K$). For any unitary representation $\rho$ of $G$ in a Hilbert apace $\mathcal{L}$, we denote by $\mathcal{L}_{\alpha}$ the subspace of $K_{\alpha}$-fixed vectors, by $P_{\alpha}$ the orthogonal projection to $\mathcal{L}_{\alpha}$. Consider operators
@@
Let $\mathcal{K}$ be a category. Denote by $\operatorname{Ob}(\mathcal{K})$ the class of objects of $\mathcal{K}$, by $\operatorname{Mor}(V,W)=\operatorname{Mor}_{\vphantom{|^{|}}\mathcal{K}}(V,W)$ the collection of all morphisms $V\to W$, by $\operatorname{End}(V)$ the semigroup of all endomorphisms of an object $V$. We assume that these semigroups have units $1_{V}$ and these units are units in the category $\mathcal{K}$, i.e., for any $\mathfrak{p}\in\operatorname{Mor}(V,W)$ we have
@@
A *-representation $(Q,\varkappa)$ of a category $\mathcal{K}$ is a functor from $\mathcal{K}$ to the category of separable Hilbert spaces and bounded linear operators. Namely, for each $V\in\operatorname{Ob}(\mathcal{K})$ we assign a Hilbert space $Q(V)$ and for each $\mathfrak{p}\in\operatorname{Mor}(V,W)$ we assign a bounded operator
@@
Consider two *-representations $(Q,\varkappa)$ and $(Q^{\prime},\varkappa^{\prime})$ of $\mathcal{K}$. An intertwiner $A:(Q,\varkappa)\to(Q^{\prime},\varkappa^{\prime})$ is a collection of bounded operators $A(V):Q(V)\to Q^{\prime}(V)$ such that for any $\mathfrak{p}\in\operatorname{Mor}(V,W)$ we have
$\varkappa_{W,V}^{\prime}(\mathfrak{p})A(V)=A(W)\varkappa_{W,V}(\mathfrak{q}).$
Two *-representations $(Q,\varkappa)$ and $(Q^{\prime},\varkappa^{\prime})$ are equivalent if there is an intertwiner $A$ such that all operators $A(V)$ are unitary.
Remark. We consider representations of categories up to equivalence. For this reason, for us, there is no difference between equivalent categories. Recall that according Bernays–Gödel version of set theory, a collection of objects of a category is a class (and, generally, not a set). By definition, all categories $\operatorname{Tra}(G,\{K_{\alpha}\})$ are small. This means that collections of objects and collections of morphisms of our categories are sets (and not ‘classes’). By aesthetic reason, sometimes we consider essentially small categories, i.e., categories equivalent to small categories. $\boxtimes$
1.4. Category of partial bijections. Recall the definition of partial bijections (see, e.g., [26], Sect. VIII.1). Let $X$, $Y$ be sets. A partial bijection $\mu:X\to Y$ is a bijection from a subset $A\subset X$ to a subset $B\subset Y$. We say that $A$ is the domain $\operatorname{dom}\mu$ of $\mu$, $B$ is the image $\operatorname{im}\mu$. A product of partial bijections $\mu:X\to Y$, $\nu:Y\to Z$ is defined in the following way: $\operatorname{dom}\nu\mu:=\mu^{-1}(\operatorname{dom}\nu\,\cap\,\operatorname{im}\mu)$ and $\nu\mu(x)=\nu(\mu(x))$ for $x\in\operatorname{dom}\nu\mu$.
We also define the pseudoinverse partial bijection, if $\mu:X\to Y$ is as above, then its pseudoinverse $\mu^{*}:X\to Y$ is the inverse bijection $B\to A$.
Clearly, we get a category with involution. Below we consider the category ${\operatorname{PB}}$ of partial bijections of finite sets. Clearly, ${\operatorname{PB}}$ is essentially small. For instance, we come to an equivalent small category, if we fix a countable set $\Omega$ and reduce a collection of objects to the set of all finite subsets of $\Omega$ (morphisms are the same). On the other hand, we can choose a set of each finite cardinality, and again we come to a small category.
@@
According Arthur Lieberman [20], any irreducible representation of the group $S_{\infty}$ is induced from an irreducible representation of some subgroup $K^{\circ}_{\alpha}$ trivial on $K_{\alpha}$.
@@
According [36] (see also exposition in [26], Sect. VIII.1-2), the following theorem holds:
@@
a) The train $\operatorname{Tra}\bigl{(}S_{\infty},\{K_{\alpha}\}\bigr{)}$ is the category whose objects are sets $\{1,\dots,\alpha\}$ and morphism are partial bijections.
b) The functor $\rho\mapsto\widetilde{\rho}_{\alpha,\beta}$ is a one-to-one correspondence between equivalence classes of unitary representations of $S_{\infty}$ and equivalence classes of *-representations of $\operatorname{Tra}\bigl{(}S_{\infty},\{K_{\alpha}\}\bigr{)}$.
So unitary representations of $S_{\infty}$ (defined up to equivalence) are in one-to-one correspondence with *-representations of the category ${\operatorname{PB}}$ since this category is equivalent to $\operatorname{Tra}\bigl{(}S_{\infty},\{K_{\alpha}\}\bigr{)}$.
Classification of *-representations of the category $\operatorname{Tra}\bigl{(}S_{\infty},\{K_{\alpha}\}\bigr{)}$ is a simple problem, see [26], Sect. VIII.1-2 (also, see below, Subsect. 2). In fact, Theorem 1.1 is equivalent to Lieberman’s description [20] of unitary representations of $S_{\infty}$.
@@ 2. Oligomorphic groups and their trains
b) For any irreducible representation $\rho$ of $G$ there is an open subgroup $H^{\circ}$ and a normal subgroup $H$ in $H^{\circ}$ of finite index such that $\rho$ is inducedl [^7] For induced representations of totally disconnected groups, see Mackey [21], [24]. Also, see below Subsect. 2. from a representation of $H^{\circ}$ trivial on $H$.
@@
$\Xi:=\Omega\setminus B,\quad G:=K(B),\quad M:=p(A)\setminus B,\quad N:=q^{-1}C\setminus B$
@@
Remark. So $\operatorname{Tra}(K)$ is realized by partial bijections and the set of morphisms is closed with respect to the involution. Therefore, it is an ’inverse category’ in the sense of [16], i.e., a categorical counterpart of inverse semigroups. $\boxtimes$
@@
b) The map $\rho\mapsto\widetilde{\rho}$ is a bijection between equivalence classes of unitary representations of $K$ and equivalence classes of *-representations of the category $\operatorname{Tra}(K)$.
@@
The proof of this statement given below in Subsect. 22 is a simple corollary of Tsankov’s Theorem 2.1, formal manipulations with ’ordered categories’ in the sense of [26], and Mackey’s technique of induced representations [21], [22], [24], [2].
@@
If $\iota(A)\subset B$, then $\iota$ is fixed by the subgroup $K(B)$. Therefore functions $F$ supported by $\mathcal{E}(A,B)$ are fixed. $\square$
@@
b) Additionally, let $(Q,\varkappa)$ be irreducible. Then a $\preceq$-minimal perfect set such that $Q(D)\neq 0$ is unique (upto $K$-equivalence).
We call this $\preceq$-minimal perfect set $D$ by the conductor of an irreducible representation.
@@
Let $(Q,\varkappa)$ be a representation of a category $\mathcal{K}$. Let $C\in\operatorname{Ob}(\mathcal{K})$. Let $\xi_{\alpha}$ be a family of vectors in $Q(C)$. For any $D\in\operatorname{Ob}(\mathcal{K})$ we consider the closed linear span $N(D)$ of all vectors $\varkappa_{\vphantom{|^{|}}D,C}(\mu)\xi_{\alpha}$, where $\mu$ ranges in $\operatorname{Mor}(C,D)$. Clearly, the family of subspaces $N(D)$ is a subrepresentation in $(Q,\varkappa)$ (we call it the cyclic span of $\xi_{\alpha}$). We say that a family $\xi_{\alpha}$ is cyclic if $N(D)=Q(D)$ for all $D$.
@@
If for a *-representation $(Q,\varkappa)$ of $\operatorname{Tra}(K)$ a subspace $Q(C)$ is cyclic, then $(Q,\varkappa)$ is uniquely determined by the representation of $\operatorname{End}(C)$ in $Q(C)$.
Proof. This is a version of the GNS-construction. Let $v_{1}$, $v_{2}\in Q(C)$, $r_{1}$, $r_{2}\in\operatorname{Mor}(C,D)$. Then
@@
Proof. Consider a sequence $A_{1}$, $A_{2}$, $\dots$ including all perfect sets in $\Omega$ defined up to the $K$-equivalence. Since $K$ is oligomorphic, it has only finite number of orbits on the space of finite subsets of a fixed cardinality. So, we can assume that numbers of elements of $A_{j}$ are (non-strictly) increasing. We take the cyclic span $(Q^{(1)},\varkappa^{(1)})$ of $R(A_{1})$ and the orthogonal complement $(R^{(1)},\rho^{(1)})$ of the cyclic span. By construction, we have $R^{(1)}(A_{1})=0$. Next, we take the cyclic span $(Q^{(2)},\varkappa^{(2)})$ of $R^{(1)}(A_{1})$ and orthogonal complement $(R^{(2)},\rho^{(2)})$ of $(Q^{(2)},\varkappa^{(2)})$ in $(R^{(1)},\rho^{(1)})$. Now we have $R^{(2)}(A_{1})=0$, $R^{(2)}(A_{2})=0$. Etc.
@@
Proof. It is sufficient to verify the statement for irreducible representations. Consider an exhaustive family of subsets $B\subset R_{1}\subset R_{2}\subset\dots\subset\Omega\setminus B$. By Lemma 2.4, for each $j$ there exists $g_{j}\in K(B)$ such that
@@
Consider the quasiregular representation $\pi$ of $K$ in $\ell^{2}(\mathcal{E}(A))$, see (2.1). For $\iota\in\mathcal{E}(A)$ denote by $e_{\iota}$ the $\delta$-function supported by $\iota$. If $\iota\in\mathcal{E}(A,B)$, then $\iota$ is fixed by all $g_{j}$. Fix $\iota(A)\not\subset B$. Let $\nu\in\mathcal{E}(A)$. There is $N$ such that $(\iota(A)\setminus B)$, $(\nu(A)\setminus B)\subset R_{N}$. For large $j\geqslant N$ we have
@@
2.7. Examples. Here we discuss an initial collection of examples of oligomorphic groups (for $1^{\lozenge}$-$4^{\lozenge}$, see [41], Sect. 5), and the corresponding categories.
$0^{\lozenge}.$ Consider the infinite symmetric group $S({\mathbb{N}})$, i.e., the group of all permutations of ${\mathbb{N}}$. All finite subsets in ${\mathbb{N}}$ are perfect, morphisms are partial bijections. This category is equivalent to the category ${\operatorname{PB}}$, see Subsect. 1.
$1^{\lozenge}.$ Consider the group $G=\operatorname{Aut}({\mathbb{Q}},<)$ of all order preserving bijections of the set ${\mathbb{Q}}$ of rational ${\mathbb{Q}}$ is a ’generic’ contable linearly ordered set. For instance, let us construct a random ordered set in the following way. We take a point $x_{1}$. With probability $1/2$ we assume $x_{2}<x_{1}$ or $x_{2}>x_{1}$. With probability $1/3$ we put $x_{3}$ to one of possible 3 positions, etc. With probability 1 we come to an ordered set equivalent to ${\mathbb{Q}}$.. Perfect sets are arbitrary finite subsets $A\subset{\mathbb{Q}}$. We come to the category $\operatorname{Tra}\bigl{(}\operatorname{Aut}({\mathbb{Q}},<)\bigr{)}$. It objects are finite subsets of ${\mathbb{Q}}$. Morphisms $A\to B$ are order preserving partial bijections. Groups of automorphisms are trivial.
This category is equivalent to the category of finite ordered sets, morphisms are order preserving partial bijections. Another equivalent category is the category, whose objects are sets $\{1,\dots,\alpha\}$, where $\alpha=0$, 1, 2, …; morphisms are order preserving partial bijections.
So, irreducible unitary representations are enumerated by the number #$\#A$ of elements of a conductor $A$. In other words, any irreducible representations of this group is realized in the spaces $\ell^{2}$ on the set of all $\alpha$-element subsets in ${\mathbb{Q}}$ for some $\alpha$.
$2^{\lozenge}.$ Let $\mathcal{R}$ be the Rado graph, see, e.g., [8]. Recall that it is a graph with a countable number of vertices satisfying the following condition (the ’Urysohn extension property’): let $A\subset B$ be finite graphs, then any embedding $A\to\mathcal{R}$ extends to an embedding $B\to\mathcal{R}$.
Consider a countable set of vertices, for each pair of vertices we draw an edge connecting them with probability $1/2$. Then we get the Rado graph with probability 1, see [10], [9].
The group $\operatorname{Aut}(\mathcal{R})$ of automorphisms of $\mathcal{R}$ acting on the set of vertices is perfect oligomorphic, arbitrary finite subset of the set of vertices is perfect. Objects of the train $\operatorname{Tra}(\operatorname{Aut}(\mathcal{R}))$ are finite subgraphs $X\subset\mathcal{R}$, a morphism $X\to Y$ is an isomorphism of an induced subgraph [^10] To avoid a misleading we use the term ’induced subgraph’: we consider some set of vertices of $X$ and all edges of $X$ connecting these vertices. $A\subset X$ to an induced subgraph $B\subset Y$.
This category is equivalent to the category, whose objects are finite graphs and morphisms are isomorphisms of induced subgraphs.
@@
$3^{\lozenge}.$ Consider a linear space $\ell(\mathbb{F})$ over a finite field $\mathbb{F}$ consisting of sequences $v=(v_{1},v_{2},\dots)$, where $v_{j}\in\mathbb{F}$ and $v_{k}=0$ for sufficiently large $k$. The group $\operatorname{GL}(\infty,\mathbb{F})$ of all linear transformations of $\ell(\mathbb{F})$ can be regarded as the group of all invertible infinite matrices over $\mathbb{F}$, which have finite number of elements in each column. The group $\operatorname{GL}(\infty,\mathbb{F})$ is perfect oligomorphic, perfect subsets are finite-dimensional linear subspaces in $\ell(\mathbb{F})$.
@@
Objects of the train are finite dimensional spaces $X$ in $\ell(\mathbb{F})$. A morphism $X\to Y$ is a partial linear bijection, i.e., an invertible operator from a subspace $A\subset X$ to a subspace $B\subset Y$. Groups of automorphisms are group of linear transformations $\operatorname{GL}(X)$.
This category is equivalent to the category of all finite-dimensional linear spaces over $\mathbb{F}$ (with same morphisms). We also can consider another equivalent category, whose objects are coordinate spaces $\mathbb{F}^{0}$, $\mathbb{F}^{1}$, $\mathbb{F}^{2}$, …with same morphisms.
So irreducible representation are enumerated by a number $n=0$, $1$, $2$, $\dots$ and a complex irreducible representation of the group $\operatorname{GL}(n,\mathbb{F})$; the picture is completely parallel to representations of $S_{\infty}$.